OpenStudy (anonymous):

How to complete the square of 4x^2+10x-3=0

OpenStudy (anonymous):

get x by it self homes

OpenStudy (campbell_st):

start by rewriting it as \[4x^2 + 10x + ? = 3 + ?\] and ? is the value of the constant in the perfect square that is added to both sides Its important to know \[ 4x^2 = (2x)^2 \] you need to know that \[(a + b)^2 = a^2 + 2ab + b^2 \] so a = 2x and 2ab = 10x so \[2 \times 2x \times b = 10x\] you need to solve for b then square the value and add it to both sides... does that make sense

OpenStudy (wolf1728):

OR try tihs 4x^2+10x =3 Divide equation by the coefficient of x^2 x^2 + 2.5x = .75 get the coefficient of x which is 2.5 divide it by 2 which is 1.25 then square it (1.25)^2 = 1.5625 and add it to both sides of the equation. x^2 + 2.5x + 1.5625 = .75 + 1.5625 Taking the square root of both sides: (x +1.25) = sqr root (2.3125) (x + 1.25) = 1.5206906326 x = 0.2706906326 sqr root (2.3125) also equals -1.5206906326 (x + 1.25) = -1.5206906326 x = -2.7706906326

OpenStudy (anonymous):

its 8

OpenStudy (wolf1728):

It is NOT 8. Putting 8 into the original equation: 4x^2+10x-3=0 4*8* + 10*8 -3 ?=0 32 + 80 -3 = 109 and so 'x' does NOT =8 the answers are: x = 0.2706906326 x = -2.7706906326